Saginaw Inc. completed its first year of operations with a pretax loss of $512,500. The tax return showed a net operating loss of $656,500, which the company wll carry forward. The $144,000 book-tax difference results from excess tax depreclation over book depreclation. Management has determined that It should record a valuation allowance equal to the net deferred tax asset. Assuming the current tax expense is zero, prepare the Journal entries to record the deferred tax provision and the valuation allowance. (If no entry is required for e transaction/event, select No Journal Entry Required In the first account field.)a. Prepare the journal entry to record the deferred tax consequences for recognition of the current year NOL before considering the valuation allowance.
Answer #1

Deferred tax asset = 656,500*34% = 223,210

Deferred tax liability = 144,000*34% = 48,960

Deferred tax benefit = (656,500-144,000)*34% = 174,250

Journal entries:

Particulars Debit(Amt) Credit(Amt)
Deferred tax asset A/c 223,210
To deferred tax benefit A/c 223,210
Deferred tax expense A/c 48,960
To deferred tax liability A/c 48,960
Deferred tax benefit A/c 174,250
To valuation allowance 174,250